OpenStudy (anonymous):

Find the vertex of the parabola. y = x2 - 16x + 73

OpenStudy (anonymous):

To find the x-value of the vertex, we will use the formula x=-b/2a. x=+16/2(1) ; x=16/2; x=8 Now that we found this x-value we pluck back into the equation and figure out why y is. 8^2-16(8)+73= 64-128+73=-64+73= y=9. Now we have found the vertex of the parabola, (8,9)

OpenStudy (anonymous):

Can you explain the reason we use that formula.